OpenStudy (anonymous):

HELP Confused:( write equation through (2,2) y-intercept 10

OpenStudy (anonymous):

y=mx+b where b is always the y-intercept Substitute: 2=m(2)+10 Solve for m. Your final solution should also be in y=mx+b form, but with only the slope and y-intercept substituted.

OpenStudy (anonymous):

okay thanks